OpenStudy (anonymous):

It takes Joan 6 times longer than jane to file the reports. Together they can file the reports in 42 minutes. How long would it take each woman to file the reports by herself?

OpenStudy (rulnick):

Jane's rate is 6x rep/min. Joan's is x rep/min. 42(6x+x)=k Jane could do it in k/6x = 49 min Joan could do it in k/x = 314 min

OpenStudy (anonymous):

Thanks!

OpenStudy (rulnick):

Correction: 314 should be 294, sorry.

OpenStudy (rulnick):

Check: 1/294 + 1/49 = 1/42.